OpenStudy (anonymous):

4cos^(2)θ+4sinθ=1 for 0°≤θ≤360°. i got 30 and 150. why wrong? @hartnn

hartnn (hartnn):

how ?

OpenStudy (anonymous):

what?

hartnn (hartnn):

i mean how u got 30 and 150 ?

OpenStudy (anonymous):

4(1-sin^(2)θ)+4sinθ=1 4-4sin^(2)θ+4sinθ=1 -4sin^(2)θ+4sinθ+3=0 sinθ=-1/2 or 1.5(rej) sinθ=-1/2 θ=30, 150...

hartnn (hartnn):

sin 30 is not -1/2 sin -30 is 1/2

hartnn (hartnn):

or, sin (360-30) = -1/2

hartnn (hartnn):

sin 30 is just +1/2

hartnn (hartnn):

same for 150

OpenStudy (anonymous):

oops

hartnn (hartnn):

so, the solutions must be 360-30 and 360-150

OpenStudy (anonymous):

thanks haha...

hartnn (hartnn):

welcome ^_^ bye, gtg

OpenStudy (anonymous):

gtg?

hartnn (hartnn):

got to go

OpenStudy (anonymous):

okok thank you a lot haha